# Linear Complexity Gibbs Sampling for Generalized Labeled Multi-Bernoulli Filtering

### v1

@@ II-A4 The GLMB Filter
All statistical information about the underlying state is contained in the filtering density–probability density of the current state conditioned on all measurements up to (and including) the current time. Given the multi-object filtering density $\boldsymbol{\pi}$ at the current time, the propagation to the next time step is given by [3]
$\boldsymbol{\pi}_{\text{+}}\!\left(\boldsymbol{X}_{\text{+}}|Z_{\text{+}}\right)$ $\propto g\left(Z_{\text{+}}|\boldsymbol{X}_{\text{+}}\right)\!\int\!\boldsymbol{\pi}\!\left(\boldsymbol{X}_{\text{+}}\right)\boldsymbol{f}\left(\!\boldsymbol{X}_{\text{+}}|\boldsymbol{X}\right)\!\delta\boldsymbol{X}.$ (3)
@@
$\boldsymbol{\pi}\left(\boldsymbol{X}\right)=\Delta\left(\boldsymbol{X}\right)\sum_{\xi\in\Xi}\left[p^{(\xi)}\right]^{\boldsymbol{X}}\sum_{I\in\mathcal{F}(\mathbb{L})}w^{\left(\xi,I\right)}\delta_{I}[\mathcal{L}\left(\boldsymbol{X}\right)],$ (4)
where $\Xi$ is some finite discrete space, $p^{\left(\xi\right)}\left(\cdot,\ell\right)$ is a probability density on $\mathbb{X}$, and $w^{\left(\xi,I\right)}$ is a non-negative weight satisfying the condition $\sum_{\left(\xi,I\right)\in\,\Xi\times\mathcal{F}(\mathbb{L})}w^{\left(\xi,I\right)}=1$. To obtain a multi-object state estimate from the GLMB density (4), we first find the most probable cardinality $n^{*}$ from the cardinality distribution, and then the highest-weighted component $(\xi^{*},I^{*})$ with cardinality $|I^{*}|=n^{*}$, see [7]. The state estimate for each object $\ell\in I^{*}$ can be taken as the mean (or mode) of $p^{(\xi^{*})}(\cdot,\ell)$. Alternatively, the entire trajectory of object $\ell\in I^{*}$ can be estimated as described in [8, 21].
@@
$\boldsymbol{\pi}\triangleq\left\{\left(p^{\left(\xi\right)},w^{\left(\xi,I\right)}\right)\right\}{}_{\left(\xi,I\right)\in\,\Xi\times\mathcal{F}(\mathbb{L})}.$ (5)
For a current GLMB filtering density of the form (5), its propagation to the next time is the GLMB [30, 8]
@@
where the parameters are
$\!\!p_{Z_{\text{+}}}^{\left(\xi,\gamma_{\text{+}}\right)}\!\left(\cdot,\ell\right)$ $=\frac{\bar{p}_{\text{+}}^{\left(\xi,\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right)\psi_{Z_{\text{+}}}^{\left(\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right)}{\bar{\psi}_{Z_{\text{+}}}^{\left(\xi,\gamma_{\text{+}}\right)}\left(\ell\right)},$ (7)
@@
$\!\!w_{Z_{\text{+}}}^{\left(\xi,\gamma_{\text{+}},I_{\text{+}}\right)}$ $\propto\delta_{\mathcal{L}(\gamma_{\text{+}})}[I_{\text{+}}]\negthinspace\!\negthinspace\sum\limits_{I\in\mathcal{F}(\mathbb{L})}\negthinspace\negthinspace\negthinspace\negthinspace w^{\left(\xi,I\right)}1_{\mathcal{\!F}(I\uplus\mathbb{B}_{\text{+}})\!\!}\left(I_{\text{+}}\right)\!w_{Z_{\text{+}}}^{\left(\xi,I,\gamma_{\text{+}}\right)}\!,$ (8)
@@
$\!\!\bar{p}_{\text{+}}^{\left(\xi,j\right)}\!\left(\cdot,\ell\right)$ =$=\begin{cases}\negthinspace\frac{\int\!f_{\text{+}}\!\left(\cdot|x,\ell\right)P_{S}\left(x,\ell\right)p^{\left(\xi\right)}\!\left(x,\ell\right)dx}{\bar{P}_{S}^{\left(\xi\right)}\left(\ell\right)},&\negthinspace\negthinspace\negthinspace\negthinspace\negthinspace\ell\in I_{\text{+}},j\!\geq 0\\ \negthinspace f_{B,\text{+}}\left(\cdot,\ell\right),&\negthinspace\negthinspace\negthinspace\negthinspace\negthinspace\!\ell\in\mathbb{B}_{\text{+}},j\!\geq 0\end{cases}$ (9)
$\!\!\bar{P}_{S}^{(\xi)}\left(\ell\right)$ $=\left\langle P_{S}\left(\cdot,\ell\right),p^{\left(\xi\right)}\left(\cdot,\ell\right)\right\rangle,$ (10)
@@
$\!\!\bar{\psi}_{Z_{\text{+}}}^{\left(\xi,\gamma_{\text{+}}\right)}\left(\ell\right)$ $=\left\langle\bar{p}_{\text{+}}^{\left(\xi,\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right),\psi_{Z_{\text{+}}}^{\left(\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right)\right\rangle,$ (11)
@@
$\!\!w_{Z_{\text{+}}}^{\left(\xi,I,\gamma_{\text{+}}\right)}$ $=1_{\Gamma_{\!\text{+}}\!}\left(\gamma_{\text{+}}\right)\prod_{\ell\in I\uplus\mathbb{B}_{\text{+}}}\eta_{Z_{\text{+}},\ell}^{(\xi,I)}(\gamma_{\text{+}}(\ell)),$ (12)
$\!\!\eta_{Z_{\text{+}},\ell}^{(\xi,I)}(j)$ =$=\!\begin{cases}1-\bar{P}_{S}^{(\xi)\!}(\ell),&\!\ell\in I,\text{}j\!<0,\\ \bar{P}_{S\!}^{(\xi)\!}(\ell)\bar{\psi}_{Z_{\text{+}}}^{(\xi,j)\!}(\ell),&\!\ell\in I,\text{ }j\!\geq 0,\\ 1-P_{B\!,\text{+}}(\ell),&\!\ell\in\mathbb{B}_{\text{+}},\text{ }j\!<0,\\ P_{B\!,\text{+}}(\ell)\bar{\psi}_{Z_{\text{+}}}^{(\xi,j)}(\ell),&\!\ell\in\mathbb{B}_{\text{+}},\text{ }j\!\geq 0.\end{cases}$ (13)
@@
Note that each component $(\xi,$$I)$ of the GLMB filtering density at time $k$ propagates a (very large) set $\left\{\left(\xi,I,\gamma_{\text{+}},I_{\text{+}}\right):I_{\text{+}}\in\mathcal{F}(\mathbb{L}_{\text{+}}),\gamma_{\text{+}}\in\Gamma_{\!\text{+}}\right\}$ of children components to the next time. Due to the terms $\delta_{\mathcal{L}(\gamma_{\text{+}})}[I_{\text{+}}]$ and $1_{\mathcal{F}(I\uplus\mathbb{B}_{\text{+}})\!}\left(I_{\text{+}}\right)$ in (8), we only need to consider components with $I_{\text{+}}\subseteq I\uplus\mathbb{B}_{\text{+}}$ and $\mathcal{L}(\gamma_{\text{+}})=I_{\text{+}}$. While this is a big reduction, in general the total number of GLMB components with non-zero weights still grows super-exponentially with time. Implementing the GLMB filter requires truncating the GLMB filtering density. Truncation by keeping the most highly weighted components minimizes the $L_{1}$ truncation error [8].
@@ II-B Gibbs Sampling for GLMB Truncation
Truncating the children of the GLMB component (indexed by) $(\xi,$$I)$ amounts to selecting the $\gamma_{\text{+}}$’s with significant $w_{Z_{\text{+}}}^{\left(\xi,I,\gamma_{\text{+}}\right)}$. For a given component $(\xi,$$I)$, let us enumerate $I=\{\ell_{1:R}\}$, $\mathbb{B}_{\text{+}\!}=\{\ell_{R\text{+}1:P}\}$, and $Z_{\text{+}}=\{z_{1:M}\}$, and abbreviate
@@
[…]
@@
$\pi(\gamma^{\prime}_{\text{+}}|\gamma_{\text{+}})=\prod\limits_{i=1}^{P}\pi_{i}(\gamma^{\prime}_{\text{+}}(\ell_{i})|\gamma^{\prime}_{\text{+}}(\ell_{1:i\texttt{-}1}),\gamma_{\text{+}}(\ell_{i\text{+}1:P})),$
where the $i$-th conditional, defined on {-$\{\text{{-}}1\mathnormal{:}M\}$, is given by
$\pi_{i}(\gamma^{\prime}_{\text{+}}(\ell_{i})|\gamma^{\prime}_{\text{+}}(\ell_{1:i\texttt{-}1}),\gamma_{\text{+}}(\ell_{i\text{+}1:P}))\propto\pi(\gamma^{\prime}_{\text{+}}(\ell_{1:i}),\gamma_{\text{+}}(\ell_{i\text{+}1:P})).$
This means, for a given $\gamma_{\text{+}}$, the next state $\gamma^{\prime}_{\text{+}}$ of the chain is generated one component after another, by sampling $\gamma^{\prime}_{\text{+}}(\ell_{i})$ from $\pi_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{1:i\texttt{-}1}),\gamma_{\text{+}}(\ell_{i\text{+}1:P}))$, $i=1,2,...,P$.
For GLMB truncation, the conditionals are the categorical distributions given by Proposition 3 of [30], which is restated in a slightly different form as follows.
@@
For each $i\in\{1$:$P\}$, let $\ell_{\bar{i}}$ denote $\ell_{1:i\texttt{-}1,i\text{+}1:P}$. Then the $i$-th conditional, defined on {-$\{\texttt{-}1\mathnormal{:}M\}$, is given by
@@
$\!\!\!\!\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$$\begin{cases}\!\eta_{i}(j),&\!j\!<\!1\\ \!\eta_{i}(j)(1-1_{\{\gamma_{\text{+}}(\ell_{\bar{i}})\}}(j)),&\!j\!\in\!\{1\mathnormal{:}M\}\end{cases}.$ (17)
@@
The above result shows that the conditionals are completely characterized by the $P\times(M+2)$ cost matrix in Fig. 1(a) (which can be pre-computed from the measurement $Z_{\text{+}}$) and the values of $\gamma_{\text{+}}$ on $\ell_{\bar{i}}$, i.e., $\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$. Specifically, the unnormalized $i$-th conditional is simply given by the $\!i$-th row after entries with (positive) indices contained in $\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$ have been zeroed (or masked) out, as illustrated in Fig. 1(b). Since evaluating $1_{\{\gamma_{\text{+}}(\ell_{\bar{i}})\}}(j)$ (and hence the mask $1-1_{\{\gamma_{\text{+}}(\ell_{\bar{i}})\}}(j)$) for each $j$ incurs an $\mathcal{O}(P)$ complexity, computing $\pi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ requires $\mathcal{O}(PM)$.
Remark 3.
@@
Selecting significant GLMB components can be performed with $\mathcal{O}(TP^{2}M)$ complexity via SGS as shown in [30]. Noting that $P$ (the number of hypothesized objects) is strongly correlated with $M$ (the number of measurements), this complexity translates roughly to a cubic complexity i.e., $\mathcal{O}(TP^{3})$ or $\mathcal{O}(TM^{3})$. Nonetheless, SGS has been extended to address multi-dimensional ranked assignment problems in multi-scan and multi-sensor GLMB filtering [8, 11, 12].
@@ III-A RGS for GLMB Truncation
Whereas SGS generates the next iterate $\gamma^{\prime}_{\text{+}}$ by traversing and updating all $P$ coordinates of the current iterate $\gamma_{\text{+}}$, RGS only selects one coordinate at random to update [41, 42]. A generic RGS implementation would incur an $\mathcal{O}(TPM)$ complexity because computing the conditionals requires $\mathcal{O}(PM)$. Further, RGS is inefficient in the sense that it generates less distinct significant samples than SGS, for the same number of iterates of the chain, leading to poorer GLMB approximations and tracking performance.

### v2

@@ II-A4 The GLMB Filter
All statistical information about the underlying state is contained in the filtering density–probability density of the current state conditioned on all measurements up to (and including) the current time. Given the multi-object filtering density $\bm{\pi}$ at the current time, the propagation to the next time step is given by [3]
$\bm{\pi}_{\text{+}}\!\left(\bm{X}_{\text{+}}|Z_{\text{+}}\right)\propto g\left(Z_{\text{+}}|\bm{X}_{\text{+}}\right)\!\int\!\bm{f}_{\text{+}}\left(\!\bm{X}_{\text{+}}|\bm{X}\right)\!\bm{\pi}\!\left(\bm{X}\right)\delta\bm{X}.$ (3)
@@
$\bm{\pi}\left(\bm{X}\right)=\Delta\left(\bm{X}\right)\sum_{\xi\in\Xi}\left[p^{(\xi)}\right]^{\bm{X}}\sum_{I\in\mathcal{F}(\mathbb{L})}w^{\left(\xi,I\right)}\delta_{I}[\mathcal{L}\left(\bm{X}\right)],$ (4)
where $\Xi$ is some finite discrete space, $p^{\left(\xi\right)}\left(\cdot,\ell\right)$ is a probability density on $\mathbb{X}$, and $w^{\left(\xi,I\right)}$ is a non-negative weight satisfying the condition $\sum_{\left(\xi,I\right)\in\,\Xi\times\mathcal{F}(\mathbb{L})}w^{\left(\xi,I\right)}=1$. To obtain a multi-object state estimate from the GLMB density (4), we first find the most probable cardinality $n^{*}$ from the cardinality distribution
$\text{Pr}(|\bm{X}|=n)\triangleq\sum_{(\xi,I)\in\Xi\times\mathcal{F}(\mathbb{L})}\delta_{n}[|I|]w^{(\xi,I)},$ (5)
and then the highest-weighted component $(\xi^{*},I^{*})$ with cardinality $|I^{*}|=n^{*}$, see [7]. The state estimate for each object $\ell\in I^{*}$ can be taken as the mean (or mode) of $p^{(\xi^{*})}(\cdot,\ell)$. Alternatively, the entire trajectory of object $\ell\in I^{*}$ can be estimated as described in [8, 17].
@@
$\bm{\pi}\triangleq\left\{\left(p^{\left(\xi\right)},w^{\left(\xi,I\right)}\right)\right\}{}_{\left(\xi,I\right)\in\,\Xi\times\mathcal{F}(\mathbb{L})}.$ (6)
Given a current GLMB filtering density of the form (6), its propagation to the next time is the GLMB [30, 8]
@@
where the new parameters are given by
$&=\frac{\bar{p}_{\text{+}}^{\left(\xi,\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right)\psi_{Z_{\text{+}}}^{\left(\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right)}{\bar{\psi}_{Z_{\text{+}}}^{\left(\xi,\gamma_{\text{+}}\right)}\left(\ell\right)},\end{aligned}$ (8)
@@
$\mathllap{\!\!w_{Z_{\text{+}}}^{\left(\xi,\gamma_{\text{+}},I_{\text{+}}\right)}}&\propto\delta_{\mathcal{L}(\gamma_{\text{+}})}[I_{\text{+}}]\negthinspace\!\negthinspace\sum\limits_{I\in\mathcal{F}(\mathbb{L})}\negthinspace\negthinspace\negthinspace\negthinspace w^{\left(\xi,I\right)}1_{\mathcal{\!F}(I\uplus\mathbb{B}_{\text{+}})\!\!}\left(I_{\text{+}}\right)\!w_{Z_{\text{+}}}^{\left(\xi,I,\gamma_{\text{+}}\right)}\!,\end{aligned}$ (9)
@@
={$&=\begin{cases}\negthinspace\frac{\int\!f_{\text{+}}\!\left(\cdot|x,\ell\right)P_{S}\left(x,\ell\right)p^{\left(\xi\right)}\!\left(x,\ell\right)dx}{\bar{P}_{S}^{\left(\xi\right)}\left(\ell\right)},&\negthinspace\negthinspace\negthinspace\negthinspace\negthinspace\ell\in I,j\!\geq 0\\ \negthinspace f_{B,\text{+}}\left(\cdot,\ell\right),&\negthinspace\negthinspace\negthinspace\negthinspace\negthinspace\!\ell\in\mathbb{B}_{\text{+}},j\!\geq 0\end{cases}\negthinspace,\end{aligned}$ (10)
[…]
@@
$&=\left\langle\bar{p}_{\text{+}}^{\left(\xi,\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right),\psi_{Z_{\text{+}}}^{\left(\gamma_{\text{+}}\left(\ell\right)\right)}\left(\cdot,\ell\right)\right\rangle,\end{aligned}$ (12)
@@
$\mathllap{\!\!w_{Z_{\text{+}}}^{\left(\xi,I,\gamma_{\text{+}}\right)}}&=1_{\Gamma_{\!\text{+}}\!}\left(\gamma_{\text{+}}\right)\prod_{\ell\in I\uplus\mathbb{B}_{\text{+}}}\eta_{Z_{\text{+}},\ell}^{(\xi,I)}(\gamma_{\text{+}}(\ell)),\end{aligned}$ (13)
={$&=\!\begin{cases}1-\bar{P}_{S}^{(\xi)\!}(\ell),&\!\ell\in I,\text{ }j\!<0\\ \bar{P}_{S\!}^{(\xi)\!}(\ell)\bar{\psi}_{Z_{\text{+}}}^{(\xi,j)\!}(\ell),&\!\ell\in I,\text{ }j\!\geq 0\\ 1-P_{B\!,\text{+}}(\ell),&\!\ell\in\mathbb{B}_{\text{+}},\text{ }j\!<0\\ P_{B\!,\text{+}}(\ell)\bar{\psi}_{Z_{\text{+}}}^{(\xi,j)}(\ell),&\!\ell\in\mathbb{B}_{\text{+}},\text{ }j\!\geq 0\end{cases}.\end{aligned}$ (14)
@@
Note that each component $(\xi,I)$ of the GLMB filtering density at time $k$ generates a (very large) set $\left\{\left(\xi,I,\gamma_{\text{+}},I_{\text{+}}\right):I_{\text{+}}\in\mathcal{F}(\mathbb{L}_{\text{+}}),\gamma_{\text{+}}\in\Gamma_{\!\text{+}}\right\}$ of children components to the next time. Due to the terms $\delta_{\mathcal{L}(\gamma_{\text{+}})}[I_{\text{+}}]$ and $1_{\mathcal{F}(I\uplus\mathbb{B}_{\text{+}})\!}\left(I_{\text{+}}\right)$ in (9), we only need to consider components with $I_{\text{+}}\subseteq I\uplus\mathbb{B}_{\text{+}}$ and $\mathcal{L}(\gamma_{\text{+}})=I_{\text{+}}$. While this is a big reduction, in general the total number of GLMB components with non-zero weights still grows super-exponentially with time. Implementing the GLMB filter requires truncating the GLMB filtering density. Truncation by keeping the most highly weighted components minimizes the $L_{1}$ truncation error [8].
@@ II-B Gibbs Sampling for GLMB Truncation
Truncating the children of the GLMB component (indexed by $(\xi,I)$) amounts to selecting the $\gamma_{\text{+}}$’s with significant $w_{Z_{\text{+}}}^{\left(\xi,I,\gamma_{\text{+}}\right)}$. For a given component $(\xi,I)$, let us enumerate $I=\{\ell_{1:R}\}$, $\mathbb{B}_{\text{+}\!}=\{\ell_{R\text{+}1:P}\}$, and $Z_{\text{+}}=\{z_{1:M}\}$, and abbreviate
@@
where $i\in\{1\mathnormal{:}P\}$, and $j\in\{\texttt{-}1\mathnormal{:}M\}$. Let $\pi$ be a (discrete) probability distribution on {-$\{\texttt{-}1\mathnormal{:}M\}^{P}$ defined by
@@
$\pi(\gamma^{\prime}_{\text{+}}|\gamma_{\text{+}})=\prod\limits_{i=1}^{P}\pi_{i}(\gamma^{\prime}_{\text{+}}(\ell_{i})|\gamma^{\prime}_{\text{+}}(\ell_{1:i\text{-}1}),\gamma_{\text{+}}(\ell_{i\text{+}1:P})),$
where the $i$-th conditional, defined on {-$\{\texttt{-}1\mathnormal{:}M\}$, is given by
$\pi_{i}(\gamma^{\prime}_{\text{+}}(\ell_{i})|\gamma^{\prime}_{\text{+}}(\ell_{1:i\text{-}1}),\gamma_{\text{+}}(\ell_{i\text{+}1:P}))\propto\pi(\gamma^{\prime}_{\text{+}}(\ell_{1:i}),\gamma_{\text{+}}(\ell_{i\text{+}1:P})).$
This means, for a given $\gamma_{\text{+}}$, the next state $\gamma^{\prime}_{\text{+}}$ of the chain is generated one component after another, by sampling $\gamma^{\prime}_{\text{+}}(\ell_{i})$ from $\pi_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{1:i\text{-}1}),\gamma_{\text{+}}(\ell_{i\text{+}1:P})),i=1,2,\cdots,P$.
For GLMB truncation, the conditionals are the categorical distributions given by [30, Proposition 3], which is restated in a slightly different form as follows.
@@
For each $i\in\{1\mathnormal{:}P\}$, let $\ell_{\bar{i}}$ denote $\ell_{1:i\text{-}1,i\text{+}1:P}$. Then the $i$-th conditional, defined on {-$\{\texttt{-}1\mathnormal{:}M\}$, is given by
@@
$\!\!\!\!\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))\triangleq\begin{cases}\eta_{i}(j),&\!j\!<\!1\\ \eta_{i}(j)(1-1_{\{\gamma_{\text{+}}(\ell_{\bar{i}})\}}(j)),&\!j\!\in\!\{1\mathnormal{:}M\}\end{cases}.$ (18)
Remark 1.
The above result shows that the conditionals are completely characterized by the $P\times(M+2)$ cost matrix in Fig. 1(a) (which can be pre-computed from the measurement $Z_{\text{+}}$) and the values of $\gamma_{\text{+}}$ on $\ell_{\bar{i}}$, i.e., $\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$. Specifically, the unnormalized $i$-th conditional is simply given by the $i$-th row after entries with (positive) indices contained in $\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$ have been zeroed (or masked) out, as illustrated in Fig. 1(b). Since evaluating $1_{\{\gamma_{\text{+}}(\ell_{\bar{i}})\}}(j)$ (and hence the mask $1-1_{\{\gamma_{\text{+}}(\ell_{\bar{i}})\}}(j)$) for each $j$ incurs an $\mathcal{O}(P)$ complexity, computing $\pi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ requires an $\mathcal{O}(PM)$ complexity.
@@
Selecting significant GLMB components can be performed with $\mathcal{O}(TP^{2}M)$ complexity via SGS as shown in [30]. Noting that $P$ (the number of hypothesized objects) is strongly correlated with $M$ (the number of measurements), this complexity translates roughly to a cubic complexity, i.e., $\mathcal{O}(TP^{3})$ or $\mathcal{O}(TM^{3})$. Nonetheless, SGS has been extended to address multi-dimensional ranked assignment problems in multi-scan and multi-sensor GLMB filtering [8, 11, 12].
@@ III-A RGS for GLMB Truncation
Whereas SGS generates the next iterate $\gamma^{\prime}_{\text{+}}$ by traversing and updating all $P$ coordinates of the current iterate $\gamma_{\text{+}}$, RGS only selects one coordinate at random to update[45, 46]. This means the transition kernel $\pi(\gamma^{\prime}_{\text{+}}|\gamma_{\text{+}})$ is given by $\frac{1}{P}\pi_{i}(\gamma^{\prime}_{\text{+}}(\ell_{i})|\gamma_{\text{+}}(\ell_{\bar{i}}))$ when $\gamma_{\text{+}}$ and $\gamma^{\prime}_{\text{+}}$ differ at most in the $i$-th coordinate and 00 otherwise. A generic RGS implementation would incur an $\mathcal{O}(TPM)$ complexity because computing the conditionals requires $\mathcal{O}(PM)$. Further, RGS is inefficient in the sense that it generates less distinct significant samples than SGS, for the same number of iterates of the chain, leading to poorer GLMB approximations and tracking performance.